OpenStudy (anonymous):

divide. (10d^9-6d^8+4d^5)/(2d^3) PLEASE HELP

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (anonymous):

@triciaal

OpenStudy (mathmate):

Hint: divide each term of the numerator by the denominator.

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

and remember that when you divide numbers or letters with the same base you subtract the powers

OpenStudy (triciaal):

example d^5 / d^2 = d^(5-2) =d^3

OpenStudy (jdoe0001):

\(\bf \cfrac{10d^9-6d^8+4d^5}{2d^3}\implies \cfrac{10d^9}{2d^3}-\cfrac{6d^8}{2d^3}+\cfrac{4d^5}{2d^3} \\ \quad \\ \cfrac{\cancel{ 10d^3 }d^6}{\cancel{ 2d^3 }}-\cfrac{\cancel{ 6d^3 }d^5}{\cancel{ 2d^3 }}+\cfrac{\cancel{ 4d^3 }d^2}{\cancel{ 2d^3 }}\)

OpenStudy (triciaal):

forgot your numbers

OpenStudy (triciaal):

@jdoe0001 forgot the numbers 10/2 etc

OpenStudy (anonymous):

wait so would it b 5d^9-3d^9+2d^5 ? @triciaal

OpenStudy (triciaal):

5d^6 - 3d^5+ 2d^2

OpenStudy (anonymous):

are u good at algebra 1B?

OpenStudy (triciaal):

hope so yes

OpenStudy (anonymous):

can u message me back @triciaal